the smallest angle of an isosceles triangle is one-fifth of one of the equal angles. the shortest side is 20m. what is the perimeter of the triangle?
Answer by cleomenius(959)

First, I found the degrees in the smallest angle.
5x + 5x + x = 180
x = 16.36 degrees, the smaller angle
So, the larger two angles are 81.8 degrees. These are the base angles.
By the law of Sins, 20/sin 16.36 = x/sin 81.8; the side came out to 70.14 cm.
2(70.14) + 20 = 160.28 cm. which is the perimeter.
Cleomenius.